OpenStudy (anonymous):

anyone here know who to use solver or goal seek in excel I need help Asap!!!! I'll give medal.

OpenStudy (anonymous):

here's what i've done so far

OpenStudy (anonymous):

okay thanks any way

OpenStudy (anonymous):

@Zale101

OpenStudy (zale101):

Vol=I*R \[Voltage=current*Resistor\] or \[(I)Current =\frac{(V) Voltage}{ (R) Resistor }\]

OpenStudy (zale101):



OpenStudy (zale101):

0.45 is the current for the first one, but how can i find the rest when you don't have the voltage or the currents for the rest :/

OpenStudy (anonymous):

i have for i1=27/130, i2=9/130, i3=9/65

OpenStudy (anonymous):

for r1=0.45 for the others i just repeated what you did with r2 and r3 which would be 0.3 and 0.225 right?

OpenStudy (zale101):

when you wrote r1, r2, and r3, are they in parallel?

OpenStudy (anonymous):

yes

OpenStudy (zale101):

i would do the same what you did right there, let me get a help from a user

OpenStudy (zale101):

@ganeshie8 I need your help here if your online

OpenStudy (anonymous):

okay cause the only thing that's getting me now is this solver or goal seeking thing

OpenStudy (zale101):

you're excel doc won't open in my laptop for some reason

OpenStudy (anonymous):

really? I'll try and resend it

OpenStudy (zale101):

yeah

OpenStudy (zale101):

yes it worked!

OpenStudy (anonymous):

ok great

OpenStudy (anonymous):

i've tried working on the solver or goal seek for 2 hours yesterday and a few hours tonight and i'm just not getting it. Im' new to that function.